# DS communications: Trends, Challenges, and Solutions

### v3

@@ II-B Link Engineering of DSOC
Designing DS optical links need to account for all link parameters and circuit characteristics. This includes the end-to-end distance between Earth and final targets, mission requirements, required capacity, atmospheric conditions, transmitter-Sun-receiver angle, etc. The received power at the target, after the receiver telescope, and before the photo-detector, can be expressed as [28]
${{P}}_{{r},{ap}}{=\ }{{P}}_{{t}}\;{{G}}_{{t}}\;{{G}}_{{r}}\;{{L}}_{{fs}}\;{{L}}_{{pt}{}}\;{{L}}_{{a}}\;{{L}}_{{c}}\;{{L}}_{{s}}\;{{\textrm{$\eta$}}}_{{t}}\;{{\textrm{$\eta$}}}_{{r}}$ (2)
@@
  • $P_{t}$ (Watt) is the transmitted power.
@@
  • $L_{fs}$ is the free space loss factor. Being inverse proportional with the link distance squared, $L_{fs}$ represents the link dominant loss factor.
  • $L_{pt}$ is the pointing loss factor. A pointing error of $\lambda/D$ radians, where $\lambda$ and $D$ are the signal wavelength and the transmitter diameter, respectively, corresponds to approximately 85% intensity reduction [53, 51].
@@
  • $L_{s}$ is the scintillation loss factor, which refers to the random optical-power fluctuations caused by atmospheric turbulence (AT) [58, 59].
@@
The various DSOC impairments reflect the importance of the link engineering in deriving the needed transmit power to meet a certain detection sensitivity and hence a service quality. It worth mentioning that the atmospheric and weather limitations are not of concern for optical space-to-space links [29, 31, 61, 62, 32]; making choosing the FSO over the RF in these links to be of a great value. This fact can be reflected in the drop of $L_{a}$, $L_{s}$, and $L_{c}$ loss terms in (2).
More investigations about potentials of the FSO technology in improving DS communications along with challenges facing its implementation are provided below.
@@ II-C DSOC Promises and Challenges
RF technology is currently dominating space communications facilities. The RF systems maturity and the cost of implementing FSO are delaying the technological migration. The nonobligatory of RF line-of-sight link availability [63] along with its tolerance to atmospheric effects and weather conditions, for example fog and snow, which can affect communications with ground stations [64] are also contributing factors to deferring the transition to FSO.
@@
Nonetheless, the RF bandwidth while being fairly limited requires licensing, and is susceptible to eavesdropping [63]. In addition, having the demand of high data rate achieved would come at the cost of increasing SWaP requirements which can be unfeasible and hence restricting RF systems performance upgrading. For instance, if we are acquiring several giga bits per second as downlink data rate from the closest distance to Mars, according to [30], we would need the Mars spacecraft to have a 5 m antenna with a transmitting power around 200 W, and a multiple of 34 m antennas in an array as the ground Earth receiver. In contrast, for optical, the Mars spacecraft would have a 1 m aperture and a transmitting power of 50 W, and the receiver telescope aperture size would be in the range of 10 m.
FSO, however, is emerging to address the shortcomings of conventional RF systems. Since FSO beams are narrow divergence, it requires significantly lower SWaP requirements than RF. Hence providing flexible systems upgrades and more secure communications [63]. Not to mention the almost limitless bandwidth with no license requirements the FSO offers [65].
Nonetheless, achieving a required performance over an optical link design increases in difficulty with its distance [29]. An optical signal power degrades proportionally to the square of the link distance. A high gain is required to make up for the very long distances loss and along with other losses discussed in previous subsection. DSOC hurdles can be compensated through a number of approaches, that come with their challenges though, which we identify below along with potential solutions.
@@
  • Increasing the power efficiency by relying on laser beams with large gain transmitter and large aperture diameters. Nonetheless, the cost of building large optical ground transmitters is considered to be an implementation concern [38]. Especially that the average power requirement for a DS link typically would be in the range of 1–20 W, and peak power is around 100–1k W [29].
@@
  • Accurately tracking the space terminals using beacons to point lasers towards them before sending data. Narrow beam requires an accurate pointing in order for the signal to be received at the destination, which is challenging considering DS distances [66]. Particularly, regarding flight systems, high accuracy of the beam pointing is needed, especially for downlink signals from DS that are typically with a beam-width in the range of 2–8 $\mu$rad. Adequate beam acquisition at destinations can be achieved, though, at the cost of computational and hardware complexities [67]. In addition, tracking and pointing beams toward mobile receivers in space require efficient design of accurate tracking algorithms.
@@
  • Relying on large gain receivers. The received power to spacecrafts would be around or less than 10-10[28]. As a consequence, that requires a high gain receiver telescopes, which can make it a challenge to obtain a functional, yet optimized design of the a spacecraft resources [29].
@@
  • Considering advanced signaling and detection techniques that maximize photon efficiency, i.e. number of bits per photon, that for instance would include, obtaining high detectors sensitivity and acquiring efficient modulation with coding schemes [68].
@@
  • Merging other technologies, including OAM for example, with FSO to get the most out of its high capacity and further enhance the quality of DS data returns [31]. More investments though need to be put forward to facilitate such a merge.
@@
There are a number of efforts being brought forth to facilitate the full implementation of the FSO technology in DS [33, 34, 30, 35, 69, 36, 70], addressing the DS FSO technology engineering challenges and maturity requirements [29, 39]. In this paper, we survey the efforts done in this direction and identify open research directions.
@@ III-A Modulation Schemes for DSOC
Conventional RF-based space communications relies commonly on binary phase shift keying (BPSK) modulation [71], where a carrier modulates data using two phases that are 180 degrees apart. BPSK has been favored in such an environment, thanks to its low bit error rate (BER) in the low signal-to-noise (SNR) ratio regime. However, for DSOC, there are other direct detection based modulation schemes of interest including on-off keying (OOK), pulse position modulation (PPM), wavelength modulation, and pulse intensity modulation. OOK and PPM are the most recommended among the methods mentioned above, thanks to their simplicity and power efficiency [31, 29]. Even though they are similar in concept, PPM is more favored than OOK for DSOC as the former is more power-efficient. On the other hand, OOK is more suitable for near-Earth links as it is cost-effective and bandwidth efficient [31].
While, OOK modulates one as a pulse and zero with the absence of a signal, M-ary PPM modulation, which is often used with photon-counting (PC) detector [29, 54, 72], works by dividing each channel symbol period into $M$ time slots. $\log_{2}M$ bits are modulated by transmitting a single pulse in one of the $M$ possible time slots of a symbol. The information in the PPM is encoded in the position of the pulse, making it robust against noise. PPM requires a high bandwidth though [73]. The good news, however, is that high bandwidth is readily offered by FSO and hence is not of much concern.
@@
where the photon energy $E=(hc)/\lambda$, $h$ is Planck's constant, $c$ is the light velocity, and $\lambda$ is the signal wavelength. $M$ is the modulation order. $P_{r}$ and $P_{n}$ are the received and noise power, respectively. $T_{slot}$ is the pulse duration. The three terms in the denominator of the second term of (3) correspond to three different behaviors when dominating. As $P_{n}\frac{2}{M-1}$, which is a noise-limited term, dominates, the capacity scales as a quadratic function of the signal power. On the other hand, $P_{r}\frac{1}{\ln(M)}$, which is the quantum-limited term, leads to a linear scaling of the capacity. Finally, the bandwidth-limited term, ${P_{r}^{2}}\frac{MT_{slot}}{\ln(M)E}$ leads to capacity saturation. Note that when the received $P_{r}$ is sufficiently high, the system capacity can be approximated as
@@
If the received power is sufficiently high, one can easily observe that it is best to use low PPM modulation order.
In Fig. 3(a) and 3(b), we show the PPM systems achieved capacity for Mars and beyond Mars distances. Here we pick the antenna diameter for transmitter and receiver as 0.22 m and 4 m [28, 38, 68], respectively. The figure shows that as the modulation order decreases, the capacity improves significantly but only up to a certain distance. That is intuitive as increasing the modulation order implies a reduction in the transmission rate and BER. As depicted in Fig. 3(b), as a signal travels a longer distance, in our settings close to 2$\times$109 km, the trend changes. As distance gets larger and larger, however, to maintain a certain BER performance, the modulation order needs to be increased. Fig. 3 indicates how crucial the increase in DS communications distance is in determining the PPM system parameters and hence performance, for a given channel model, detector, and other system parameters including receiver antenna diameter and transmission power, as discussed in [70] and a number of references therein. The impact of combining the PPM modulation technique with coding is also of crucial importance as we will discuss.
@@ III-B Coding for DSOC Links
Channel coding is a very important block in the transmission chain to enhance performance and ensure reliability. Applying ideal error control coding (ECC) may lead to meeting a link capacity limit, maximum possible achievable capacity. However, such ideal codes are hard to construct in practice. Therefore, it is important to consider the code efficiency, which refers to the difference between the capacity limit and the error code threshold at a specific BER, when designing communications links [76].
@@
In the early missions to DS, around sixties, a decision was needed on whether to rely over a coding method or the other, namely block codes or convolutional codes. During the design of Mariner '69 Mars mission, for example, block codes were the best options, given that developed sequential decoding algorithms were not that efficient for DS. Such a choice led to a poor data rate, though. Thanks to the development of a more efficient sequential decoding algorithm, convolutional decoding methods, for example Turbo codes, have been adopted to various communications channels, including DS [77], starting from Pioneer 9 mission [78], which was designed after Mariner '69 but lunched earlier. Such a change offered a coding gain, which is the power gain over the uncoded case, of 3 dB [79] compared to that of 2.2 dB [78] in the case of block code used by Mariner '69.
@@
Reed-Solomon (RS) codes are commonly used with convolutional codes for DS application as they are efficient in correcting complex and bursty errors [41, 54, 29, 80], and have in the order of 2.5 to 3 dB efficiency [53]. As a reference, an uncoded PPM-based system would have a code efficiency in the order of 5 dB [73]. Nonetheless, the most recommended and efficient DS coding scheme is the serially concatenated PPM (SCPPM[81, 82, 73]. This coding technique has a code efficiencies between 0.5 and 1 dB [53, 73]. Fig. 4 shows the BER of SCPPM versus average signal power measured in signal photons detected per modulation order and denoted by $P_{av}$, compared to the capacity limit, Reed-Solomon PPM (RSPPM), and uncoded PPM. The figure confirms the superiority of the SCPPM over RSPPM, as it is clearly demonstrating the lower coding efficiency and higher gain SCPPM achieves over RSPPM.
@@
Recently, there have also been some suggested codes in the field. For example, Divsalar et al. in [83] show that coded optical wavelength division multiple access (WDMA) with PPM can be a good candidate for DS. Using the coding gain metric, WDMA has a 7 dB gain for a 0.8 coding rate, and a 12 dB coding gain for a coding rate of 1/6 at a frame error rate (FER) of 10-6 [83].
@@
Since fixed-rate codes like low-density parity-check (LDPC) and Turbo codes are not efficient for DS communications as they fail to adapt to the changes in communications link due to time variability, Liang et al. in [84] proposed the use of rate-less codes such as raptor codes, and hence achieve a better performance than LDPC and Turbo codes. The authors propose to further combine spinal and polar coding to achieve even better performance.
@@
Further improvement can be offered through the use of efficient detection method as we will elaborate next.
@@ III-C Detection Techniques for DSOC
Advances in detection technology play an important role in providing reliable DSOC. As indicated in Section II, having improved sensitivity and efficient detection are regarded as means to overcome faint DS signals issues and design limitations aspects. Detection receivers can either be coherent or non-coherent. Coherent receivers can have a higher spectral and dimensional information efficiency in one state system, examples of which are heterodyne and homodyne receivers. Dimensional information efficiency, according to [85], is defined as the efficiency of information transfer per a unit dimension or optical communications degree of freedom. Even though coherent receivers suffer from limited photon efficiency, they tend to be more practical when very high data rates are required. Non-coherent receivers, on the other hand, are energy efficient but come at the cost of an exponentially decreasing dimensional efficiency [85]. Rendering to their energy efficiency, non-coherent receivers which are alternatively called direct detectors are commonly used in DS [53, 86]. There are different types of direct detectors among which are avalanche photodiode detector (APD), and PIN (p-type, intrinsic, and n-type) diode, and PC detector.
@@
Detectors/ Receivers Type
Implementation
Sensitivity (PPB)
Data rate (Mbps)
Advantages
Disadvantages
InGaAs
Direct PIN
Cubesats for DSOC systems
$163$
Relatively high temperature operation
@@
Relatively high received power requirement, not a good option as a ground receiver due to atmospheric losses
GmAPD
APD
@@
Mars laser communications demonstration (MLCD) project, imaging system for DSOC Project 2022
2
$34.9$
No read out noise penalty
Limited bandwidth
SNSPD
PC
Lunar laser communications demonstration (LLCD) demonstrations ,DSOC Project 2022
0.5
$781$
High detection efficiency ($90\%$)
@@
Requirement for cooling down to 2–4 K, high photon rate detection (multiple of Gbps) can be hard to attain
SQ
Coherent Homodyne
Near-Earth and DS applications
1.5
156
Energy efficient
Limited bandwidth
SPC
Heterodyne
EDRS, DS communications (suggested)
10
1800
@@
Robust to severe background noise, near-Sun communications
High bandwidth requirement
PSA
Pre-amplified
DS application demonstrations
1
10500
Noise-free amplification
@@
High performance in limited spectral efficiency range
@@
Due to their high sensitivity compared to other conventional direct detection methods, PC detectors are gaining great attention and are recommended along with PPM modulation to be used in the future DSOC. These detectors can count a single photon incident on them. PC detectors can be viewed as an extension of APD detectors with infinite gain in which a digital output signal is generated for each detected photon [86]. They can offer up to 20 dB gain along with PPM over APD and pre-amplified direct detectors [29], and sensitivities of a few photons per bit [74]. The PC detection process, though, is affected by noise sources which are $i)$ background noise that comes from the diffused energy from the sky, planets, or stars in the field of view of the receiver, $ii)$ dark counts electrons generated by the detector, and $iii)$ receiver thermal noise [28]. In addition to noise, the losses that degrade the PC detection efficiency are $i)$ Blocking loss which results from the fact that a detector gets blind after a detection event missing all incident photons until reset. Blocking loss limits a detector to count at most one photon per reset time or dead time [60]. $ii)$ Jitter loss which is the random delay elapses from the incident of a photon on a detector to the time an electrical output pulse is generated in response to that photon. A significant loss occurs if the standard deviation of the jitter is in the order of the slot width of the pulse [87].
The choice of the detection and modulation schemes is essential in mitigating other losses resulting from link impairments. Detectors that handle a small duty cycle in the range of a few nanoseconds, for example, are resilient to long DSOC channel losses [88]. An example of a detector that can provide such a requirement is the superconducting nanowire single photon detector (SNSPD[89], which is a commercialized highest-performing PC detector in the frequency range of 30–30,000 THz [90]. This detector can support slot widths down to 200 ps with detection efficiencies of 50–78% [89]. There are some other examples of detectors and receivers that are significant in DS communications, such as indium gallium arsenide (InGaAs[91], Geiger-mode avalanche photodiode (GmAPD[92, 93, 94], SNSPD [95, 96, 97, 98, 90], single-quadrature (SQ[99], European data relay system (EDRS) with single photon coherent (SPC) heterodyne detection [100, 101], and phase-sensitive optical amplifiers (PSA[74].

### v4

@@ II-B Link Engineering of DSOC
Designing DS optical links needs to account for all link parameters and communicating entities’ characteristics. This includes the end-to-end distance between Earth and final targets, mission requirements, required capacity, atmospheric conditions, transmitter-Sun-receiver angle, etc. The received power at the target after the receiver telescope and before the photo-detector can be expressed as [28]:
${{P}}_{tar}{=\ }{{P}}_{{t}}\;{{G}}_{{t}}\;{{G}}_{{r}}\;{{L}}_{{fs}}\;{{L}}_{{pt}{}}\;{{L}}_{{a}}\;{{L}}_{{c}}\;{{L}}_{{s}}\;{{\textrm{$\eta$}}}_{{t}}\;{{\textrm{$\eta$}}}_{{r}},$ (2)
@@
  • $P_{t}$ is the transmitted power in Watts.
@@
  • $L_{fs}$ is the free space loss factor. Being inversely proportional to the link distance squared, $L_{fs}$ represents the link dominant loss factor.
  • $L_{pt}$ is the pointing loss factor. A pointing error of $\lambda/D_{t}$ radians, where $\lambda$ is the signal wavelength and $D_{t}$ is the transmitter diameter, corresponds to approximately 85% intensity reduction [53, 51].
@@
  • $L_{s}$ is the scintillation loss factor, which refers to the random optical power fluctuations caused by the atmospheric turbulence (AT) [58, 59].
@@
The various DSOC impairments reflect the importance of the link engineering in deriving the needed transmit power to meet a certain detection sensitivity and hence a service quality. It is worth mentioning that the atmospheric and weather limitations are not of concern for optical space-to-space links [29, 31, 61, 62, 32], making the choice of the FSO over the RF in these links of great value. This fact can be reflected in the drop of $L_{a}$, $L_{s}$, and $L_{c}$ loss terms in (2).
More investigations about the potentials of the FSO technology in improving DS communications along with challenges facing its implementation are provided below.
@@ II-C DSOC Promises and Challenges
RF technology is currently dominating space communications facilities. The RF systems’ maturity and the cost of implementing FSO are delaying the technological migration. The nonobligatory of RF line-of-sight link availability [63] along with its tolerance to atmospheric effects and weather conditions [64] are also contributing factors to deferring the transition to FSO.
@@
Nonetheless, the fairly limited RF bandwidth requires licensing. Furthermore, RF communications are susceptible to eavesdropping [63]. In addition, having the demand for high-data rates comes at the cost of increasing SWaP requirements which can be unfeasible and hence restrict RF system performance upgrades. For instance, if we are acquiring several gigabits per second as downlink data rate from the closest distance to Mars, according to [30], we would need the Mars spacecraft to have a 5 m antenna with a transmitting power of around 200 W, and a multiple of 34 m antennas in an array as the ground Earth receiver. In contrast, for optical, the Mars spacecraft would have a 1 m aperture and a transmitting power of 50 W, and the receiver telescope aperture size would be in the range of 10 m.
FSO, on the other hand, is emerging to address the shortcomings of conventional RF systems. FSO beams, as they expand thousands of times slower than RF [65], demand significantly lower SWaP allowing more flexible system upgrades and secure communications [63]. Not to mention the almost limitless bandwidth with no license requirements the FSO offers [66]. Nonetheless, achieving high performance over DS distances is challenging [29]. A high gain is required to make up for the very long distances loss along with other losses discussed in the previous subsection. DSOC implementation hurdles can be compensated through several approaches, that come with their challenges, as we discuss below.
@@
  • Increasing the power efficiency by relying on laser beams with large-gain transmitters and large-aperture diameters. Nonetheless, the cost of building large optical ground transmitters is considered to be an implementation concern [38]. Especially since the average power requirement for a DS link typically would be in the range of 1–20 W and peak power would be ranging from 100 to 1k W [29].
@@
  • Accurately tracking the space terminals using beacons to point lasers towards them before sending data. A narrow beam requires accurate pointing in order for the signal to be received at the destination, which is challenging considering DS distances [67]. Particularly, regarding flight systems, high accuracy of the beam pointing is needed, especially for downlink signals from DS that are typically with a beam-width in the range of 2–8 $\mu$rad. Adequate beam acquisition at destinations can be achieved at the cost of computational and hardware complexities though [68]. In addition, tracking and pointing beams toward mobile receivers in space require efficient design of accurate tracking algorithms.
@@
  • Relying on large-gain receivers. The received power to spacecrafts would be around or less than 10-10[28]. As a consequence, that requires high gain receiver telescopes, which can make it a challenge to obtain a functional yet optimized design of the spacecraft resources [29].
@@
  • Considering advanced signaling and detection techniques that maximize photon efficiency, i.e., the number of bits per photon. Achieving this would also include getting high-detection sensitivity and acquiring efficient modulation and coding schemes [69].
@@
  • Merging FSO with other technologies, including OAM for example, to get the most out of its high capacity and further enhance the quality of DS data returns [31]. More investments though need to be put forward to facilitate such a merge.
@@
There are several efforts being brought forth to facilitate the full implementation of the FSO technology in DS [33, 34, 30, 35, 70, 36, 71] and address the DS FSO technology engineering challenges and maturity requirements [29, 39]. In this paper, we investigate the efforts made in DSOC and identify open research directions.
@@ III-A Modulation Schemes for DSOC
Conventional RF-based space communications rely commonly on binary phase shift keying (BPSK) modulation [72], where a carrier modulates data using two phases that are 180 degrees apart. BPSK has been favored in such an environment, thanks to its low bit error rate (BER) in the low signal-to-noise (SNR) ratio regime. However, for DSOC, there are other direct detection-based modulation schemes of interest including on-off keying (OOK), pulse position modulation (PPM), wavelength modulation, and pulse intensity modulation. OOK and PPM are the most recommended among the methods mentioned above, thanks to their simplicity and power efficiency in photon-starved regimes [31, 29].
As PPM is more power efficient, it is more preferred than OOK for DSOC. On the other hand, OOK is more suitable for near-Earth links as it is bandwidth-efficient and cost-effective [31]. While, OOK modulates one as a pulse and zero in the absence of a signal, M-ary PPM modulation, which is often used with photon-counting (PC) detector [29, 54, 73], works by dividing each channel symbol period into $M$ time slots. $\log_{2}M$ bits are modulated by transmitting a single pulse in one of the $M$ possible time slots of a symbol. The information in the PPM is encoded in the position of the pulse, making it robust against noise. PPM requires a high bandwidth though [74]. The good news, however, is that the high bandwidth is readily offered by FSO and hence it is not of much concern.
@@
where the photon energy $E=(hc)/\lambda$, $h$ is Planck’s constant, $c$ is the light velocity, and $\lambda$ is the signal wavelength. $M$ is the modulation order. $P_{r}$ and $P_{n}$ are the received and noise power, respectively. $T_{slot}$ is the pulse duration. The three terms in the denominator of the second term of (3) correspond to three different behaviors when dominating. As $P_{n}\frac{2}{M-1}$, which is a noise-limited term, dominates, the capacity scales as a quadratic function of the signal power. On the other hand, $P_{r}\frac{1}{\ln(M)}$, which is the quantum-limited term, leads to a linear scaling of the capacity. Finally, the bandwidth-limited term, ${P_{r}^{2}}\frac{MT_{slot}}{\ln(M)E}$ leads to capacity saturation. Note that when the received $P_{r}$ is sufficiently high, the system capacity can be approximated as
@@
If the received power is sufficiently high, one can easily observe that it is best to use a low PPM modulation order.
In Fig. 3(a) and 3(b), we show the PPM systems achieved capacity for Mars and beyond Mars distances. Here we pick the antenna diameter for the transmitter as 0.22 m and the receiver as 4 m [28, 38, 69]. The figure shows that as the modulation order decreases, the capacity improves significantly but only up to a certain distance. That is intuitive as increasing the modulation order implies a reduction in the transmission rate and BER. As depicted in Fig. 3(b), as a signal travels a longer distance, the trend changes. I.e., as the distance gets larger and larger, to maintain a certain BER performance, the modulation order may need to be increased. Fig. 3 indicates how crucial the increase in DS communications distance is in determining the PPM system parameters and hence performance, as discussed in [71] and several references therein. The impact of combining the PPM modulation technique with coding is also of crucial importance as we will discuss.
@@ III-B Coding for DSOC Links
Channel coding is a very important block in the transmission chain to enhance performance and ensure reliability. Applying ideal error correction coding (ECC) may lead to meeting a link capacity limit, the maximum possible achievable capacity. However, such ideal codes are hard to construct in practice. Therefore, it is important to consider code efficiency, which refers to the difference between the capacity limit and the error code threshold at a specific BER, when designing communications links [77].
@@
In the early missions to DS, in the 1960s, a decision was needed on whether to rely on a coding method or the other, namely block codes or convolutional codes. During the design of Mariner ’69 Mars mission, for example, block codes were the best options given that the developed sequential decoding algorithms were not that efficient for DS. Such a choice led to a poor data rate, though. Thanks to the development of a more efficient sequential decoding algorithm, convolutional decoding methods, for example, Turbo codes have been adopted for various communications channels including DS [78], starting from Pioneer 9 mission [79] which was designed after Mariner ’69 but launched earlier. Such a change offered a coding gain, which is the power gain over the uncoded case, of 3 dB [80] compared to that of 2.2 dB [79] in the case of the block code used by Mariner ’69.
@@
Reed-Solomon (RS) codes are commonly used with convolutional codes for DS applications as they are efficient in correcting complex and bursty errors [41, 54, 29, 81], and have in the order of 2.5 to 3 dB efficiency [53]. As a reference, an uncoded PPM-based system would have a code efficiency in the order of 5 dB [74]. Nonetheless, the most recommended and efficient DS coding scheme is the serially concatenated PPM (SCPPM[82, 83, 74]. This coding technique has a code efficiency between 0.5 and 1 dB [53, 74]. Fig. 4 shows the BER of SCPPM versus average signal power measured in single photons detected per modulation order and denoted by $P_{av}$, compared to the capacity limit, Reed-Solomon PPM (RSPPM), and uncoded PPM. The figure confirms the superiority of SCPPM over RSPPM, as it clearly demonstrates lower coding efficiency and higher gain.
@@
Recently, there have been other codes suggested for DS. For example, Divsalar et al. [84] show that coded optical wavelength division multiple access (WDMA) with PPM can be a good candidate. Using the coding gain metric, WDMA has a 7 dB gain for a 0.8 coding rate and a 12 dB coding gain for a coding rate of 1/6 at a frame error rate (FER) of 10-6 [84].
@@
Since fixed-rate codes like low-density parity check (LDPC) and Turbo codes are not efficient for DS communications as they fail to adapt to changes incurring in communications links due to time variability, Liang et al. [85] proposed the use of rate-less codes such as raptor codes, and hence achieve a better performance than LDPC and Turbo codes. The authors propose to further combine spinal and polar coding to achieve even better performance.
@@
Further improvement can be offered through the use of efficient detection methods as we will elaborate on next.
@@ III-C Detection Techniques for DSOC
Advances in detection technology play an important role in providing reliable DSOC. As indicated in Section II, having improved sensitivity and efficient detection are regarded as means to overcome faint DS signals issues and design limitations. Detection receivers can either be coherent or non-coherent. Coherent receivers can have a higher spectral and dimensional information efficiency in one state system, examples of which are heterodyne and homodyne receivers. Dimensional information efficiency, according to [86], is defined as the efficiency of information transfer per unit dimension or optical communications degree of freedom. Even though coherent receivers suffer from limited photon efficiency, they tend to be more practical when very high data rates are required. Non-coherent receivers, on the other hand, are energy efficient but come at the cost of exponentially decreasing dimensional efficiency [86]. Rendering to their energy efficiency, non-coherent receivers which are alternatively called direct detectors are commonly used in DS [53, 87]. There are different types of direct detectors among which are avalanche photodiode detectors (APD), PIN (p-type, intrinsic, and n-type) diodes, and PC detectors.
@@
Detectors/Receivers Type Implementation Sensitivity (PPB) Data rate (Mbps) Advantages Disadvantages
@@
InGaAs Direct PIN Cubesats for DSOC systems - $163$ Relatively high-temperature operation Relatively high received power requirement, not a good option as a ground receiver due to atmospheric losses
@@
GmAPD APD Mars laser communications demonstration (MLCD) project, imaging system for DSOC Project 2 $34.9$ No read out noise penalty Limited bandwidth
@@
SNSPD PC Lunar laser communications demonstration (LLCD) demonstrations, DSOC Project 0.5 $781$ High-detection efficiency ($98\%$) Requirement for cooling down to 2–4 K, high photon rate detection (multiple of Gbps) can be hard to attain
SQ Coherent Homodyne Near-Earth and DS applications 1.5 156 Energy efficient Limited bandwidth
@@
SPC Heterodyne EDRS, DS communications (suggested) 10 1800 Robust to severe background noise, near-Sun communications High-bandwidth requirement
@@
PSA Pre-amplified DS application demonstrations 1 10500 Noise-free amplification High performance in a limited spectral efficiency range
@@
Due to their high sensitivity compared to other conventional direct detection methods, PC detectors are gaining great attention and are recommended along with PPM modulation to be used in the future DSOC. These detectors can count a single photon incident on them. PC detectors can be viewed as an extension of APD detectors with infinite gain in which a digital output signal is generated for each detected photon [87]. They can offer up to 20 dB gain along with PPM over APD and pre-amplified direct detectors [29], and sensitivities of a few photons per bit [75]. The PC detection process, though, is affected by noise sources which are $i)$ background noise that comes from the diffused energy from the sky, planets, or stars in the field of view of the receiver, $ii)$ dark counts of electrons generated by the detector, and $iii)$ receiver thermal noise [28]. In addition to noise, the losses that degrade the PC detection efficiency are $i)$ blocking loss which results from the fact that a detector gets blind after a detection event missing all incident photons until reset. Blocking loss limits a detector to counting at most one photon per reset time or dead time [60], and $ii)$ jitter loss which is the random delay that elapses from the incident of a photon on a detector to the time an electrical output pulse is generated in response to that photon. A significant loss occurs if the standard deviation of the jitter is in the order of the slot width of the pulse [88]. The choice of the detection and modulation schemes is essential in mitigating other losses resulting from link impairments. Detectors that handle a small duty cycle in the range of a few nanoseconds, for example, are resilient to long DSOC channel losses [89]. An example of a detector that can provide such a requirement is the superconducting nanowire single photon detector (SNSPD[90, 91, 92, 93, 94, 95], which is the highest-performing commercialized PC detector in the frequency range of 30–30,000 THz [95]. This detector can support slot widths down to 200 ps, timing jitter less than 10 ps [96], and up to 98% detection efficiency [97, 90]. There are some other examples of detectors and receivers that are significant in DS communications, such as indium gallium arsenide (InGaAs[98], Geiger-mode avalanche photodiode (GmAPD[99, 100, 101], single quadrature (SQ[102], European data relay system (EDRS) with single photon coherent (SPC) heterodyne detection [103, 104], and phase-sensitive optical amplifiers (PSA[75].